OpenStudy (anonymous):

Convert to the given unit. 40 km to m A. 4000 m B. 40,000 m C. 400,000 m D. 4,000,000 m

OpenStudy (jojokiw3):

1 Km = 1000 m. So therefore...40 km = ?

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} km&m \\\hline\\ 1&1000\\ 40&x \end{array}\implies \cfrac{1}{40}=\cfrac{1000}{x}\)

OpenStudy (anonymous):

a right?

OpenStudy (jojokiw3):

What's 40 x 1000?

sammixboo (sammixboo):

\(\sf \Large \color{magenta}{\frac{1}{40}}=\color{aqua}{\frac{1000}{?}}\) We know \(\sf \large \color{magenta}{1 ~x~1000 ~=~1000}\) so what is \(\sf \large \color{aqua}{40~x~1000}\)

OpenStudy (anonymous):

4000 right?

OpenStudy (jojokiw3):

There are forty 1000 meters. How many meters are there?

OpenStudy (anonymous):

i said 4000

OpenStudy (jojokiw3):

Okay. If there are four 1000 meters, it would be 4000. so 4 x 1000 = 4000. BUUT. 1 km = 1000 m. 40 km = ?

OpenStudy (anonymous):

wait no it is 40,000

sammixboo (sammixboo):

\(\sf \large \color{magenta}{4~x~1000 =~4000}\) We are trying to find out what is \(\sf \large \color{magenta}{40~x~1000 }\)

sammixboo (sammixboo):

Yes

OpenStudy (jojokiw3):

Mhm.

OpenStudy (anonymous):

ok thx ppl